Choose h and k such that the system below has (a) no solution, (b) a unique solution, and (c)
many solutions.

x1 + 3x2 = 2

3x1+ hx2 = k.
1
Expert's answer
2020-12-01T03:01:31-0500

Let us consider the system


"\begin{cases} 
x_1 + 3x_2 = 2\\
3x_1+ hx_2 = k
\end{cases}" (*)


Let us find the determinant "\Delta=\left|\begin{array}{cc}1 & 3 \\ 3 & h\end{array}\right|=h-9" .

(a) If "\Delta=0", then "h=9", and the system (*) is equivalent to the system "\begin{cases} 
x_1 + 3x_2 = 2\\
x_1+ 3x_2 = \frac{k}{3}
\end{cases}" .

If "\frac{k}{3}\ne 2", that is "k\ne 6", then the system (*) has no solution.



(b) If "\Delta\ne 0", that is "h\ne 9", then for any "k" the system (8) has a unique solution.



(c) If "\Delta=0", that is "h=9", and "k=6", then the system (*) is equivalent to the system


"\begin{cases} 
x_1 + 3x_2 = 2\\
x_1+ 3x_2 =2
\end{cases}" , and therefore, the system (*) has infinitely many solutions.